OpenStudy (anonymous):

Formula for surface area, volume, circumference, volume of composite figures.

OpenStudy (anonymous):

hey @camerondoherty fan me and message me please

OpenStudy (anonymous):

Could use a little help thanks.

OpenStudy (anonymous):

hey cam I told u that pic would get u fanboys xD wait a sec skater im getting the formulas

OpenStudy (anonymous):

sorray okay ill help

OpenStudy (anonymous):

Thanks (:

OpenStudy (camerondoherty):

The formula for Volume is: \[L * W * H = V\] L = Legnth W = width H = Heght V = Volume The formula for Circumference is: \[C = \pi d\] \[\pi \]=pi C = circumference d = diameter

OpenStudy (camerondoherty):

More coming up lol

OpenStudy (anonymous):

Thank you so much

OpenStudy (camerondoherty):

Well for the other formulas there are many... it depends on what shape or object you are finding the volume for...

OpenStudy (anonymous):

I'm not so good with composite figures...

OpenStudy (anonymous):

wait cam aren't the formulas different depending on the figures?

OpenStudy (camerondoherty):

yea thts wut i said

OpenStudy (camerondoherty):

not all of them though volume and circumference are for circles and rectangles, cubes, etc

OpenStudy (anonymous):

what about composite figures? Those are the ones that get me confused

OpenStudy (anonymous):

Helpp?